# Waterloo, Ontario, N2L 2Y5, Canada

### v1

@@ 6 Discussion
In the shell geometry of section 4, the geometry is locally AdS3 for large radii, i.e., $r>R$, and we found the $u<d$ criterion continues to control the scattering process as long as the relevant RT surfaces only probe this region. So, according to our interpretation above, the matter shell only affects the key entanglement structure of the corresponding boundary state at low energies or at large separations. Now, conical defects with $M=-1/n^{2}$ are $\mathbb{Z}_{n}$ orbifolds, and following Balasubramanian:2014sra ; Balasubramanian:2016xho , one may “ungauge” the discrete $\mathbb{Z}_{n}$ gauge symmetry to produce a covering theory where the $O(c)$ internal degrees of freedom in the original CFT are now split up into $n$ boundary subregions in covering CFT, with $O(c/n)$ internal degrees of freedom. This “entwinement” trick was used in Caminiti:2024ctd to convert $u<d$ to a statement about internal degrees of freedom. So, we are claiming that for the shell geometry, the internal entanglement structure which Caminiti:2024ctd associated with holographic scattering is unchanged in the UV regime, and disrupted only in the IR. We comment that the disruption of entanglement by bulk matter is reminiscent of the discussion of BTZ black holes in Caminiti:2024ctd ; in this case, the entanglement structure is disrupted by correlations to the thermofield double degrees of freedom. We also observe that it would be interesting to understand how the entwinement trick extends to stationary examples such as spinning defects (as discussed in appendix B) or to dynamic configurations with multiple defects Balasubramanian:1999zv .
@@
First, we observe that the proof of the connected wedge theorem in May:2019odp ; May:2021nrl involves a lower bound on $I(\hat{\mathcal{V}}_{1}:\hat{\mathcal{V}}_{2})$ in terms of the focusing [^35] More generally, quantum focusing Bousso:2015mna . properties of the lightsheets defining the scattering region. In pure AdS3, the expansion of these lightsheets vanishes, leading to a saturation of the bound. We expect one could leverage this saturation in order to prove the CWT converse in pure AdS3 (as we did in section 2.2, albeit via direct calculations). Now, in the presence of matter, the expansion of the lightsheets can be nonvanishing, and we wonder whether a careful analysis can be used to upgrade $u<d$ to a matter-dependent statement; for example, $d-\tilde{u}>m$, where $m>0$ depends on the matter content of the geometry. As we have seen in the shell geometry of section 4, in some cases, the $u$ surface does not exist, so here $\tilde{u}$ represents a possible generalization of the $u$ surface which does always exist.
@@
To propose a possible definition of $\tilde{u}$, let us revisit the key conclusions of section 4.3. In this section, we found that there is no equivalence between holographic scattering and a competition among nonminimal RT surfaces in the shell geometry. However, we did find that holographic scattering implies certain relations among nonminimal RT candidates when $\theta<\Delta\phi^{*}$. (For $\theta>\Delta\phi^{*}$, analogous results were obtained using the numerical methods of appendix C, albeit on a case-by-case basis.) In particular, if the $u$ surface exists, we found holographic scattering implies $u<d$; otherwise, if the $i$ surface exists, it implies $i<d$; and otherwise, $o<d$. These statements are nontrivial because in some cases, $o$ is the minimal RT candidate, and so holographic scattering need only imply $o<d$, and yet we find the additional relations $i<d$ and $u<d$. As mentioned in section 4.3, inspired by the upcoming work ap ; ap2 , we can summarize these statements by writing $u_{simple}<d$, where we define $u_{simple}$ as the unique candidate in $\{o,u,i\}$ which is contained in the simple wedge of the boundary interval with opening angle $\theta+x$. From this result, a natural proposal is to define $\tilde{u}$ with reference to the $\theta+x$ simple-wedge, such that in the shell geometry, $\tilde{u}=u_{simple}$. Then, an idea for future work is to understand if focusing implies the existence of some $m$ such that $d-\tilde{u}<m$ matches the scattering condition.
@@
A promising indication in this direction is that focusing arguments can directly explain the $u_{simple}<d$ effect just described, as we have learned from upcoming work ap ; ap2 . This is a sign that focusing arguments indeed have more to teach us about the relationship between holographic scattering and nonminimal surfaces. Another interesting topic for future work is to investigate the possibility of using similar focusing arguments to prove the aforementioned conjecture about the redundancy of the $\gamma_{\hat{\mathcal{R}}_{1}}$ assumption in the $u<d$ converse of section 3.
@@
Stepping back, we note that since a general converse to the CWT is not known, efforts have been made by other authors to introduce modifications of the CWT that admit a converse. In this paper, in the spirit of Caminiti:2024ctd , we aimed to find extra conditions, in addition to connectedness of the entanglement wedge, which are sufficient to imply scattering. However, another possible approach would be to find a new definition of the scattering region (typically, larger than the traditional scattering region) such that connectedness of the entanglement wedge alone is sufficient to imply “scattering”; for example, see Leutheusser:2024yvf ; ap .
@@
To elaborate on our analysis, we showed that the only cases where the scattering threshold matches the boundary between connected and disconnected entanglement wedges in pure AdS3 is with $n=2$, $n=3$ with $\theta\leq 2\pi/n$, and $n>3$ at the $\mathbb{Z}_{n}$-symmetric point: $\theta=2\pi/n$. It would be interesting to understand what is so special about the entanglement structure at the $\mathbb{Z}_{n}$-symmetric point, such that having a connected wedge is always enough to imply scattering. It would also be interesting to compare the scattering curves with phase diagrams for more exotic RT candidates, such as RT candidates which do not obey the standard homology constraint, or alternatively, to the behavior of proposed multipartite entanglement measures in holography such as Balasubramanian:2014hda ; Bao:2018gck ; Bao:2019zqc ; Akers:2019gcv ; Dutta:2019gen ; DeWolfe:2020vjp ; Gadde:2022cqi ; Agon:2022efa ; Hayden:2011ag ; Ju:2024kuc ; Balasubramanian:2024ysu . Alternatively, rather than looking for a precise match to the scattering curve, one could first try using focusing arguments to show that scattering implies relations among nonminimal surfaces for the $n$-to-$n$ case, as in the $u_{simple}<d$ discussion above.
@@ Appendix A Geodesics and cuts in pure AdS3
In section 5, we also make use of the notion of future lightcone cuts Engelhardt:2016wgb . As stated in the main text, the future lightcone cut of a bulk point $p$ is defined as the intersection of $p$’s future lightcone (that is, the boundary of the future of $p$) with the conformal boundary of the manifold. That is, an event at $p$ in the bulk can only influence boundary points on or to the future of this cut. In the context of AdS3 in global coordinates (86), the lightcone cut simply defines a closed curve on the two-dimensional boundary. Using eq. (90), one straightforwardly obtains the lightcone cut of the bulk point at $({r_{p}},{t_{p}},{\phi_{p}})$ as Engelhardt:2016wgb
@@ Appendix B Example: spinning defect
Spinning defects in AdS3 (e.g., see Miskovic:2009uz ; Compere:2018aar ; Edelstein:2011vu ; Briceno:2021dpi ; Li:2024rma ) are described by the following stationary but not static geometry:
@@
where 1<M<0-1<M<0, $r>0$, $\phi\sim\phi+2\pi$, and $t\in\mathbb{R}$. We focus our attention on $|M|\geq|J|$, which is motivated by the recent KSW criteria Kontsevich:2021dmb ; Witten:2021nzp for the admissibility of complex metrics in the gravitational path integral Basile:2023ycy . Further, unitarity of the dual CFT implies $|J|\leq 1+M$ raeymaekers2011 ; raeymaekers2010 , so we will restrict ourselves to the interval $|J|\leq\min(|M|,1-|M|)$. We note the same metric yields the spinning BTZ black hole when $M\geq 0$ Banados:1992wn ; Banados:1992gq , while we recover pure AdS3 with M=1,J=0M=-1,J=0.
@@
After the following change of coordinates, [^37] Introducing null coordinates $u_{0}=t-\phi$, $v_{0}=t+\phi$, $u=\tau-\varphi$, $v=\tau+\varphi$, the coordinate change on the boundary reads simply $u=(b_{+}-b_{-})u_{0}$, $v=(b_{+}+b_{-})v_{0}$.
@@
A subtlety here is that this identification leads to a spacetime with closed timelike curves. To see why, recall that pure AdS3 comes with the coordinate range $0\leq\rho^{2}<\infty$. Under eq. (95), this corresponds to $b_{-}^{2}\leq r^{2}<\infty$ Li:2024rma , and in particular, $r^{2}$ can be negative. Now consider the closed curve $\phi(s)=2\pi s$, $s\in[0,1]$, with $t$ and $r$ held fixed, or in global AdS3 coordinates
@@
To explicitly check that $u<d$ controls scattering, we first find spacelike geodesics in the spinning defect geometry (92) in order to evaluate the holographic entanglement entropy. These can be obtained by mapping pure AdS3 RT surfaces (87) to the spinning defect geometry, using eq. (95). For a spacelike geodesic which extends between two points $(t_{i},\phi_{i})$ and $(t_{f}=t_{i},\phi_{f})$ on the boundary, [^38] In principle, geodesics in the spinning defect geometry can wind around $r=0$ multiple times Martinez:2019nor , but we will not be interested in these cases. we obtain
@@
where $\epsilon$ sets the UV, i.e., we choose the cutoff surface close to the boundary at $r=\frac{1}{\epsilon}$, and recall $b_{+}\pm b_{-}=\sqrt{|M|\pm J}$. The length of the second candidate is given by eq. (103) upon substituting $\Delta\phi\mapsto 2\pi-\Delta\phi$. With $J=0$, we recover the expressions for the static conical defect Caminiti:2024ctd .
@@
Using eq. (103), we can construct the entanglement wedge phase diagram in the spinning defect geometry with M=0.5,J=0.499M=-0.5,J=0.499. The phase boundaries are shown in black in figure 13 below. Compared to the $J=0$ case, there is a larger portion of the phase diagram for which the entanglement wedge is connected. However, this effect is very small – on the order $\Delta x\sim 10^{-3}$, and so the phase diagram closely resembles figure 8 of Caminiti:2024ctd .
@@
We determine the scattering condition below and find a perfect match to (105), as expected from section 3 and illustrated in figure 13. This is therefore an example where $u<d$ implies holographic scattering. Note that in this case, $\Delta\phi_{CTC}\approx 1.1\pi<2\pi$, so there are choices of $\theta$ and $x$ for which the $u$ surface crosses into the CTC region (namely, when $\theta+x>\Delta\phi_{CTC}$). As pointed out in Caminiti:2024ctd , there exist regimes where $u<d$ controls scattering even when $o$ is the minimal candidate, and hence in these cases scattering implies a relation among non-minimal RT candidates.
@@ Appendix C Details for shell geometry
We use the Israel junction conditions Israel:1966rt (e.g., see also Misner:1973prb ) to determine the stress tensor of our shell, such that Einstein’s equations are obeyed. First, we evaluate the extrinsic curvature of the $r=R$ surface for the interior and exterior geometries, which is simply given by

### v3

@@ 6 Discussion
In the shell geometry of section 4, the geometry is locally AdS3 for large radii, i.e., $r>R$, and we found the $u<d$ criterion continues to control the scattering process as long as the relevant RT surfaces only probe this region. So, according to our interpretation above, the matter shell only affects the key entanglement structure of the corresponding boundary state at low energies or at large separations. Now, conical defects with $M=-1/n^{2}$ are $\mathbb{Z}_{n}$ orbifolds, and following Balasubramanian et al. (2015, 2016), one may “ungauge” the discrete $\mathbb{Z}_{n}$ gauge symmetry to produce a covering theory where the $O(c)$ internal degrees of freedom in the original CFT are now split up into $n$ boundary subregions in covering CFT, with $O(c/n)$ internal degrees of freedom. This “entwinement” trick was used in Caminiti et al. (2024) to convert $u<d$ to a statement about internal degrees of freedom. So, we are claiming that for the shell geometry, the internal entanglement structure which Caminiti et al. (2024) associated with holographic scattering is unchanged in the UV regime, and disrupted only in the IR. We comment that the disruption of entanglement by bulk matter is reminiscent of the discussion of BTZ black holes in Caminiti et al. (2024); in this case, the entanglement structure is disrupted by correlations to the thermofield double degrees of freedom. We also observe that it would be interesting to understand how the entwinement trick extends to stationary examples such as spinning defects (as discussed in appendix B) or to dynamic configurations with multiple defects Balasubramanian and Ross (2000).
@@
First, we observe that the proof of the connected wedge theorem in May et al. (2020); May (2021) involves a lower bound on $I(\hat{\mathcal{V}}_{1}:\hat{\mathcal{V}}_{2})$ in terms of the focusing [^35] More generally, quantum focusing Bousso et al. (2016). properties of the lightsheets defining the scattering region. In pure AdS3, the expansion of these lightsheets vanishes, leading to a saturation of the bound. We expect one could leverage this saturation in order to prove the CWT converse in pure AdS3 (as we did in section 2.2, albeit via direct calculations). Now, in the presence of matter, the expansion of the lightsheets can be nonvanishing, and we wonder whether a careful analysis can be used to upgrade $u<d$ to a matter-dependent statement; for example, $d-\tilde{u}>m$, where $m>0$ depends on the matter content of the geometry. As we have seen in the shell geometry of section 4, in some cases, the $u$ surface does not exist, so here $\tilde{u}$ represents a possible generalization of the $u$ surface which does always exist.
@@
To propose a possible definition of $\tilde{u}$, let us revisit the key conclusions of section 4.3. In this section, we found that there is no equivalence between holographic scattering and a competition among nonminimal RT surfaces in the shell geometry. However, we did find that holographic scattering implies certain relations among nonminimal RT candidates when $\theta<\Delta\phi^{*}$. (For $\theta>\Delta\phi^{*}$, analogous results were obtained using the numerical methods of appendix C, albeit on a case-by-case basis.) In particular, if the $u$ surface exists, we found holographic scattering implies $u<d$; otherwise, if the $i$ surface exists, it implies $i<d$; and otherwise, $o<d$. These statements are nontrivial because in some cases, $o$ is the minimal RT candidate, and so holographic scattering need only imply $o<d$, and yet we find the additional relations $i<d$ and $u<d$. As mentioned in section 4.3, inspired by the upcoming work Lima et al. (2025b, a), we can summarize these statements by writing $u_{simple}<d$, where we define $u_{simple}$ as the unique candidate in $\{o,u,i\}$ which is contained in the simple wedge of the boundary interval with opening angle $\theta+x$. From this result, a natural proposal is to define $\tilde{u}$ with reference to the $\theta+x$ simple-wedge, such that in the shell geometry, $\tilde{u}=u_{simple}$. Then, an idea for future work is to understand if focusing implies the existence of some $m$ such that $d-\tilde{u}<m$ matches the scattering condition.
@@
A promising indication in this direction is that focusing arguments can directly explain the $u_{simple}<d$ effect just described, as we have learned from upcoming work Lima et al. (2025b, a). This is a sign that focusing arguments indeed have more to teach us about the relationship between holographic scattering and nonminimal surfaces. Another interesting topic for future work is to investigate the possibility of using similar focusing arguments to prove the aforementioned conjecture about the redundancy of the $\gamma_{\hat{\mathcal{R}}_{1}}$ assumption in the $u<d$ converse of section 3.
@@
Stepping back, we note that since a general converse to the CWT is not known, efforts have been made by other authors to introduce modifications of the CWT that admit a converse. In this paper, in the spirit of Caminiti et al. (2024), we aimed to find extra conditions, in addition to connectedness of the entanglement wedge, which are sufficient to imply scattering. However, another possible approach would be to find a new definition of the scattering region (typically, larger than the traditional scattering region) such that connectedness of the entanglement wedge alone is sufficient to imply “scattering”; for example, see Leutheusser and Liu (2025); Lima et al. (2025b).
@@
To elaborate on our analysis, we showed that the only cases where the scattering threshold matches the boundary between connected and disconnected entanglement wedges in pure AdS3 is with $n=2$, $n=3$ with $\theta\leq 2\pi/n$, and $n>3$ at the $\mathbb{Z}_{n}$-symmetric point: $\theta=2\pi/n$. It would be interesting to understand what is so special about the entanglement structure at the $\mathbb{Z}_{n}$-symmetric point, such that having a connected wedge is always enough to imply scattering. It would also be interesting to compare the scattering curves with phase diagrams for more exotic RT candidates, such as RT candidates which do not obey the standard homology constraint, or alternatively, to the behavior of proposed multipartite entanglement measures in holography such as Balasubramanian et al. (2014b); Bao and Halpern (2019); Bao and Cheng (2019); Akers and Rath (2020); Dutta and Faulkner (2021); DeWolfe et al. (2020); Gadde et al. (2022); Agón et al. (2023); Hayden et al. (2013); Ju et al. (2025); Balasubramanian et al. (2025). Alternatively, rather than looking for a precise match to the scattering curve, one could first try using focusing arguments to show that scattering implies relations among nonminimal surfaces for the $n$-to-$n$ case, as in the $u_{simple}<d$ discussion above.
@@ Appendix A Geodesics and cuts in pure AdS3
In section 5, we also make use of the notion of future lightcone cuts Engelhardt and Horowitz (2017). As stated in the main text, the future lightcone cut of a bulk point $p$ is defined as the intersection of $p$’s future lightcone (that is, the boundary of the future of $p$) with the conformal boundary of the manifold. That is, an event at $p$ in the bulk can only influence boundary points on or to the future of this cut. In the context of AdS3 in global coordinates (86), the lightcone cut simply defines a closed curve on the two-dimensional boundary. Using eq. (90), one straightforwardly obtains the lightcone cut of the bulk point at $({r_{p}},{t_{p}},{\phi_{p}})$ as Engelhardt and Horowitz (2017)
@@ Appendix B Example: spinning defect
Spinning defects in AdS3 (e.g., see Miskovic and Zanelli (2009); Compère and Fiorucci (2018); Edelstein et al. (2011); Briceño et al. (2021); Li (2024)) are described by the following stationary but not static geometry:
@@
where 1<M<0-1<M<0, $r>0$, $\phi\sim\phi+2\pi$, and $t\in\mathbb{R}$. We focus our attention on $|M|\geq|J|$, which is motivated by the recent KSW criteria Kontsevich and Segal (2021); Witten (2021) for the admissibility of complex metrics in the gravitational path integral Basile et al. (2023). Further, unitarity of the dual CFT implies $|J|\leq 1+M$ Raeymaekers et al. (2011, 2010), so we will restrict ourselves to the interval $|J|\leq\min(|M|,1-|M|)$. We note the same metric yields the spinning BTZ black hole when $M\geq 0$ Banados et al. (1992, 1993), while we recover pure AdS3 with M=1,J=0M=-1,J=0.
@@
After the following change of coordinates, adapted from Li (2024), [^37] Introducing null coordinates $u_{0}=t-\phi$, $v_{0}=t+\phi$, $u=\tau-\varphi$, $v=\tau+\varphi$, the coordinate change on the boundary reads simply $u=(b_{+}-b_{-})u_{0}$, $v=(b_{+}+b_{-})v_{0}$.
@@
A subtlety here is that this identification leads to a spacetime with closed timelike curves. To see why, recall that pure AdS3 comes with the coordinate range $0\leq\rho^{2}<\infty$. Under eq. (95), this corresponds to $b_{-}^{2}\leq r^{2}<\infty$, and in particular, $r^{2}$ can be negative. Now consider the closed curve $\phi(s)=2\pi s$, $s\in[0,1]$, with $t$ and $r$ held fixed, or in global AdS3 coordinates
@@
To explicitly check that $u<d$ controls scattering, we first find spacelike geodesics in the spinning defect geometry (92) in order to evaluate the holographic entanglement entropy. These can be obtained by mapping pure AdS3 RT surfaces (87) to the spinning defect geometry, using eq. (95). For a spacelike geodesic which extends between two points $(t_{i},\phi_{i})$ and $(t_{f}=t_{i},\phi_{f})$ on the boundary, [^38] In principle, geodesics in the spinning defect geometry can wind around $r=0$ multiple times Martínez et al. (2019), but we will not be interested in these cases. we obtain
@@
where $\epsilon$ sets the UV, i.e., we choose the cutoff surface close to the boundary at $r=\frac{1}{\epsilon}$, and recall $b_{+}\pm b_{-}=\sqrt{|M|\pm J}$. The length of the second candidate is given by eq. (103) upon substituting $\Delta\phi\mapsto 2\pi-\Delta\phi$. With $J=0$, we recover the expressions for the static conical defect Caminiti et al. (2024).
@@
Using eq. (103), we can construct the entanglement wedge phase diagram in the spinning defect geometry with M=0.5,J=0.499M=-0.5,J=0.499. The phase boundaries are shown in black in figure 13 below. Compared to the $J=0$ case, there is a larger portion of the phase diagram for which the entanglement wedge is connected. However, this effect is very small – on the order $\Delta x\sim 10^{-3}$, and so the phase diagram closely resembles figure 8 of Caminiti et al. (2024).
@@
We determine the scattering condition below and find a perfect match to (105), as expected from section 3 and illustrated in figure 13. This is therefore an example where $u<d$ implies holographic scattering. Note that in this case, $\Delta\phi_{CTC}\approx 1.1\pi<2\pi$, so there are choices of $\theta$ and $x$ for which the $u$ surface crosses into the CTC region (namely, when $\theta+x>\Delta\phi_{CTC}$). As pointed out in Caminiti et al. (2024), there exist regimes where $u<d$ controls scattering even when $o$ is the minimal candidate, and hence in these cases scattering implies a relation among non-minimal RT candidates.
@@ Appendix C Details for shell geometry
We use the Israel junction conditions Israel (1966) (e.g., see also Misner et al. (1973)) to determine the stress tensor of our shell, such that Einstein’s equations are obeyed. First, we evaluate the extrinsic curvature of the $r=R$ surface for the interior and exterior geometries, which is simply given by